OpenStudy (philcua):

The formula d = rt gives the distance d traveled in time t at rate r. A bicyclist rides at a constant rate of 22 kmph. How many kilometers will he travel in 2.5 h? km=

OpenStudy (philcua):

@Abhisar @AriPotta

OpenStudy (philcua):

Can you help

OpenStudy (aripotta):

so the rate, r, is 22 and the time, t, would be 2.5. plug them into the formula to solve for d d = rt d = 22(2.5)

OpenStudy (philcua):

which would be 55 @AriPotta

OpenStudy (aripotta):

yea that should be your answer i have to something real quick. i'll be back to help you in like 10 mins

OpenStudy (philcua):

Ok

OpenStudy (philcua):

@geerky42

OpenStudy (anonymous):

This one?

OpenStudy (philcua):

Yeah I will post the next question now

OpenStudy (philcua):

Solve the problem. The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

Apply that formula \(d=rt\). We are given t and d. So \(\large r=\frac{d}{t}\)

OpenStudy (philcua):

which would be 41/2.3 ?

OpenStudy (anonymous):

Yes?

OpenStudy (philcua):

So it would would be 17.82

OpenStudy (anonymous):

Exactly

OpenStudy (philcua):

so c but it is 17.83

OpenStudy (aripotta):

that's because you round up, not down. the answer is 17.826... so that would be rounded to 17.83

OpenStudy (philcua):

Last one The formula P = ns gives the formula for the perimeter of a regular polygon with n sides and side length s. What is the length of each side of a regular octagon with a perimeter of 54.4 cm? A regular octagon has 8 equal sides. cm =

OpenStudy (aripotta):

use the formula to solve for s p = ns s = p/n s = 54.4/8

OpenStudy (philcua):

Which would be 6.8

OpenStudy (aripotta):

ye

OpenStudy (philcua):

Ok my scores are coming in as we speak

OpenStudy (philcua):

100000000000% thank you guys yoour the best

OpenStudy (aripotta):

yay :)

OpenStudy (philcua):

Ok gotta go cya

OpenStudy (anonymous):

Bye